# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning

### v1

@@ II-B2 Masked Image Modeling
In the field of RS, RingMo [1] pioneered the foundation model framework based on MIM. Unlike random masking techniques used in natural scenes, RingMo developed the PIMask strategy tailored to the complex backgrounds and small targets characteristic of RS data, focusing on preserving information about small targets. By integrating ViT and Swin Transformer, RingMo enhanced the accuracy of various downstream RS tasks, laying a foundation for developing large models in RS scenes.
II-B3 The combination of Contrastive Learning &\&& Masked Image Modeling
@@
Despite the success of models like ViT and RingMo in general and satellite remote sensing tasks, their performance in aerial remote sensing remains limited. Currently, there is no dedicated foundation model for aerial remote sensing. To address this gap, we propose RingMo-Aerial, a model designed specifically for aerial remote sensing tasks, and demonstrate its effectiveness in common applications.
III Problem Description
The unique imaging characteristics of ARS photography stem from the inherent properties of aerial perspective. Specifically, this perspective manifests as a size disparity within the same image due to the inclined angle of capture. Objects that are closer to the ARS (towards the bottom of the image) appear larger, while those further away (towards the top) appear smaller, as illustrated in Figure 5.
To elaborate, as depicted in Figure 6, when a ARS captures two objects of equal actual size at inclined angles of $\theta_{1}$ and $\theta_{2}$, the resulting pixel dimensions and their ratio can be mathematically expressed. For illustration, consider two identical cars on a highway, both with a height of L. The vertical component of the cars in the camera’s view can be expressed as $L\times sin(\theta_{1})$ and $L\times sin(\theta_{2})$. Given the drone’s flight height is $H$, the focal length of the drone’s camera is $f$, and the horizontal distances between the two cars and the drone are $d_{1}$ and $d_{2}$ respectively, the slant distances between the cars and the drone are calculated as $D_{1}=\sqrt{H^{2}+d_{1}^{2}}$ and $D_{1}=\sqrt{H^{2}+d_{2}^{2}}$. According to the principle of perspective projection, the height of the objects in pixels in the image is calculated as follows:
$h_{1}=\frac{f\cdot L\cdot\sin(\theta_{1})}{D_{1}}$ (1)
$h_{2}=\frac{f\cdot L\cdot\sin(\theta_{2})}{D_{2}}$ (2)
The ratio of the heights of the pixels occupied by the cars is:
$\frac{h_{1}}{h_{2}}=\frac{D_{2}}{D_{1}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}=\frac{\sqrt{H^{2}+d_{2}^{2}}}{\sqrt{H^{2}+d_{1}^{2}}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}$ (3)
Then the proportion of pixels occupied by car 1 and car 2 can be estimated as:
$\frac{p_{1}}{p_{2}}=\frac{h_{1}^{2}}{h_{2}^{2}}$ (4)
@@ III Problem Description
Figure 5: Similar targets with significant size differences in the same picture. In the Vision Transformer model, objects of different sizes occupy different numbers of tokens.The image is from VisDrone dataset.
Figure 6: The difference in shooting angles causes the final pixel sizes of similar targets to be different.
@@
Considering that in the visual transformer model, the number of pixels occupied by an object is related to the number of tokens it occupies, as 5. Therefore, it is obviously useful to adopt a method to balance the number of tokens occupied by similar objects of different sizes in an image.
@@
Figure 7: The main structure of RingMo-Aerial.
IV Methodology
IV-A Method Overview
The proposed RingMo-Aerial model serves as a foundational framework with broad applicability. Its implementation involves three main stages: model design, pretraining, and fine-tuning for downstream tasks. In Section IV-B, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase. To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown as Figure 7. In Section IV-C, we combine the commonly used masked image modeling (MIM) pretraining with an additional contrastive learning (CL) method based on affine transformations during the pretraining stage. This hybrid pretraining approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown as Figure 7. Furthermore, in Section IV-D, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase. The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks, shown as Figure 10.
IV-B Model Architecture
RingMo-Aerial’s design is based on Swin Transformer. The architecture of Swin Transformer consists of four stages, each comprising either a linear encoder or Patch Merging, along with an even number of Swin Transformer Blocks. Each block includes Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA). These attention mechanisms effectively compute attention maps within the windows, extracting spatial feature information between pixels while reducing the computational complexity of the model.
@@ IV-B Model Architecture
To address the aforementioned challenges, we propose the overall architecture of RingMo-Aerial. As shown in Figure 7(a), the overall architecture of RingMo-Aerial is based on the Swin Transformer. Considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
@@
Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module. To better understand the FE-MSA mechanism, it is essential to first recognize the limitations of the existing window attention mechanisms. These mechanisms often struggle to balance the semantic information in feature vectors of objects of different sizes. For instance, as illustrated in Figure 8, larger objects may occupy $4\times 4$ tokens, while smaller distant objects may occupy only $3\times 3$ tokens. The FE-MSA mechanism introduces frequency domain processing to the existing window attention, indirectly expanding the receptive field for smaller objects.
As shown in Figure 7(b), the RingMo-Aerial Block structure is illustrated, including two window attention mechanisms optimized by the FE-MSA mechanism. Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block. Taking the first stage as an example, the input feature vector $z^{l}$ adopts the same slicing parameters as the standard Swin-Transformer, resulting in four-channel slices $C=\{c_{1},c_{2},c_{3},c_{4}\}$. After linear normalization, it enters the FE-MSA module. For each patch with a dimension of $C$, a $7\times 7$ convolution kernel is used for depthwise convolution to obtain the feature map $Fz^{(l-1)}=\text{Concat}\{Fc_{1},Fc_{2},Fc_{3},Fc_{4}\}$. Then, W-MSA or SW-MSA is used to compute the attention map. After the FE-MSA, the edge and texture features of small objects are further enhanced, improving the model’s ability to perceive both local and global features. The computation process for each token patch through the RingMo-Aerial block is as follows:
@@
The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, enhancing the model’s performance in complex aerial scenes.
IV-C Pre-train
Figure 9: Positive samples for contrastive learning are obtained after affine transformation, cropping and resizing of the image.
For pretrain, we construct a large-scale pre-training dataset. Our comprehensive pre-training dataset covers a wide range of aerial scenes, including high-altitude, low-altitude, and near-ground perspectives. It spans diverse geographical regions such as urban, rural, and wilderness areas. The dataset includes various targets, such as roads, vehicles, and pedestrians, and captures different seasons and time periods. Notably, it exhibits multi-source, multi-temporal, and multi-instance characteristics. Prior to training, we crop the entire dataset into 640x640 or 448x448 resolution slices, all captured using optical three-channel imaging. In total, the sliced dataset comprises approximately 5.4 million images, as Table I.
TABLE I: The Number of Post-Cropping Images Per Dataset.
Datasets Images
AU-AIR [42] 328230
DAC-SDC [43] 73740
DarkTrack2021 [44] 223606
DroneCrowd [45] 42960
Okutama-Action [46] 530874
Seadronessee [47] 913986
StanfordDroneDataset [48] 580223
UAVDark135 [49] 124980
UZH-FPV [50] 184458
VTUAV [51] 1576800
WebUAV3M [52] 376903
DTB70 [53] 47331
ERA [54] 39720
ITCVD [55] 27680
UAV123 [56] 11161
UAVHuman [57] 22549
UAV-ROD [58] 42088
UCLA Aerial Event dataset [59] 27370
Internet collection 228316
Total 5402975
During the pre-training phase, we adopt a combination of two techniques: Masked Image Modeling (MIM) and Contrastive Learning (CL). For MIM training, we use the standard SimMIM pre-training framework. The primary goal of this approach is to enhance the model’s ability to capture pixel-level semantic associations, thus initially equipping the model with feature extraction capabilities. However, as analyzed previously, solely using the MIM pre-training method for ARS foundational model training is insufficient for the model to learn the intrinsic object characteristics specific to ARS vision. Therefore, building upon the MIM pre-training, we propose a contrastive learning framework based on affine transformations. This involves generating positive and negative image pairs through affine transformations, which serve as the basis for the contrastive learning pre-training of the model.
@@ IV-C Pre-train
Among them, $\alpha$ is the transformation coefficient, and the value range is randomly between 0.05 and 0.35. On this basis, the perspective transformation matrix can be calculated. The perspective transformation matrix is defined as a $3\texttimes 3$ matrix in the form of:
@@
The solution process of this transformation matrix is:
@@
Through the aforementioned affine transformation process, we obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$. The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs. By calculating the contrastive loss, we guide the model to learn the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
IV-D Downstream Fine-tuning
Figure 10: The proposed ARS-Adapter.
Figure 11: Qualitative results for the UAVID dataset when using different methods.
In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks. Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model. The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, as 10. The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking. In a fashion akin to AdaptFormer [60], the introduced ARS-Adapter module operates in parallel with the model. The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer. Therefore, the forward propagation is:
[…]
Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling. The Adapter in parallel with FFN introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
[…]
V Experiments
TABLE II: Performance Comparison on AID and RESISC-45 Datasets
Model AID (OA) RESISC-45 (OA)
(TR=20%/50%) (TR=10%/20%)
GASSL [61] 93.55/95.92 90.86/93.06
SeCo [33] 93.47/95.99 89.64/92.91
SatMAE [62] 95.02/96.94 91.72/94.10
RingMo [1] 96.90/98.34 94.25/95.67
RVSA [63] 97.03/98.50 93.93/95.69
DINO-MC [64] 95.16/97.09 -
TOV [65] 95.16/97.09 90.97/93.79
SSL4EO [66] 91.06/94.74 87.60/91.27
CMID [67] 96.11/97.79 91.84/94.07
CACo [68] 90.88/95.05 88.28/91.94
CROMA [69] 96.44/97.58 92.63/95.04
SatLas [70] 94.96/97.38 92.16/94.70
GFM [71] 95.47/97.09 -
Scale-MAE [72] 96.44/97.58 92.63/95.04
SkySense [2] 97.68/98.60 94.85/96.32
RingMo-Aerial 95.81/96.46 92.28/95.65
TABLE III: Compare our model with other methods on LEVIR-CD and CDD dataset
Methods LEVIR-CD CDD
F1 IoU F1 IoU
FC-Siam-Di [73] 86.31 83.31 70.61 54.57
FC-Siam-Conc [73] 83.69 76.77 75.11 60.14
IFNet [74] 88.13 82.93 90.30 71.91
STANet [75] 87.30 77.40 84.12 72.22
BIT [76] 89.31 89.37 88.90 80.01
SNUNet [77] 88.16 87.17 83.89 72.11
ChangeFormer [78] 90.40 88.80 89.83 81.53
RingMo [1] 89.53 82.38 90.90 84.09
RingMo-Aerial (Ours) 92.71 87.10 94.62 90.09
TABLE IV: Comp. on UAVID

### v3

@@ 2.2.2 Masked Image Modeling
In RS, RingMo [1] pioneered the foundation model framework based on MIM. Unlike random masking techniques used in natural scenes, RingMo developed the PIMask strategy tailored to the complex backgrounds and small targets characteristic of RS data, focusing on preserving information about small targets. By integrating ViT and Swin Transformer, RingMo enhanced the accuracy of various downstream RS tasks, laying a foundation for developing large models in RS scenes.
2.2.3 The combination of Contrastive Learning &\&& Masked Image Modeling
@@
Despite the success of models like ViT and RingMo in general and satellite remote sensing tasks, their performance in ARS remains limited. Currently, there is no dedicated foundation model for ARS. To address this gap, we propose RingMo-Aerial, a model designed specifically for ARS tasks, and demonstrate its effectiveness in typical applications.
3 Problem Description
@@
Figure 5: Similar targets with significant size differences in the same picture. In the Vision Transformer model, objects of different sizes occupy different numbers of tokens. The image is from the VisDrone dataset.
The unique imaging properties of ARS photography stem from the inherent properties of aerial perspective. Specifically, this perspective manifests itself in size differences within the same image due to the tilted angle of the shot. As shown in Fig.5, similar objects at different distances from the ARS are rendered at different sizes in the image.
Figure 6: The same target will be represented differently in the image if its distance from the lens and imaging angle are different. As the horizontal distance increases, the number of pixels occupied by the height content increases and then decreases, and the height content occupies the most pixels at the horizontal distance $l=\sqrt{H(H-h)}$.
As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\alpha$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera. At different horizontal distances, the imaging size corresponding to the same target height differs. Specifically, when the target is directly beneath the camera, its height is not captured in the image. As the horizontal distance increases, the number of pixels occupied by the target’s height initially increases and then decreases. By considering the distances from the target’s lowest and highest points to the camera, given by $x=\sqrt{H^{2}+l^{2}}$ and $y=\sqrt{(H-h)^{2}+l^{2}}$ respectively, we can derive the relationship between the viewing angle $\alpha$ and the horizontal distance $l$ between the target and the camera:
Using the cosine rule:
$x^{2}+y^{2}-2xy\cos\alpha=h^{2}$ (1)
Substituting the values of $x$ and $y$, The final cosine formula becomes:
$\cos\alpha=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
By setting $a=H$ and $b=H-h$, we have the target function is:
$f(l)=\cos\alpha=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
Taking the derivative of $f(l)$:
$f^{\prime}(l)=\frac{2l(a^{2}+l^{2})(b^{2}+l^{2})-l(ab+l^{2})(a^{2}+b^{2}+2l^{2})}{(a^{2}+l^{2})^{3/2}(b^{2}+l^{2})^{3/2}}$ (4)
Solving $f^{\prime}(l)=0$ results in the equation:
$2(a^{2}+l^{2})(b^{2}+l^{2})=(ab+l^{2})(a^{2}+b^{2}+2l^{2})$ (5)
Simplifying further:
$2a^{2}b^{2}+a^{2}l^{2}+b^{2}l^{2}=a^{3}b+ab^{3}+2abl^{2}$ (6)
Combining terms:
$l^{2}(a-b)^{2}=ab(a-b)^{2}$ (7)
Since $a=H$ and $b=H-h$, and $a>b$, it follows that $a-b=h>0$.
When $f^{\prime}(l)=0$, we obtain:
$l^{2}=ab$ (8)
Therefore:
$l=\sqrt{ab}=\sqrt{H(H-h)}$ (9)
This completes the derivation of the maximum value $\alpha$ and the corresponding minimum value $\cos\alpha$. From Eq. (2), when $l=0$, $\alpha$ is equal to 0. From Eqs.(4) and (9), $\alpha$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\alpha$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\alpha$ is again equal to 0. Assuming that the optical axis of the camera is always at the center of the target viewpoint $\alpha$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\alpha$ of the target. This indicates that if the same target is at different distances from the lens and imaging angles, it will have different representations in the image.
Figure 7: RingMo-Aerial’s two-stage pre-training principles Capturing Pixel-level Semantic Associations and Learning Aerial Visual Physical Characteristics, and the architecture of the FE-MSA module.
@@
In the visual transformer model, the number of pixels occupied by an object is related to the number of tokens it occupies, as Fig.5. Therefore, it is helpful to adopt a method to balance the number of tokens occupied by similar objects of different sizes in an image.
@@
4 Methodology
4.1 Method Overview
The proposed RingMo-Aerial model serves as a foundational framework with broad applicability. Its implementation involves three main stages: model design, pre-training, and fine-tuning for downstream tasks. In Section 4.2, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase. To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown in Fig.7. In Section 4.3, we combine the commonly used masked image modeling (MIM) pre-training with additional contrastive learning (CL) methods based on affine transformations during the pre-training stage. This hybrid pre-training approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown in Fig.7. Furthermore, in Section 4.4, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase. The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks.
4.2 Model Architecture
RingMo-Aerial’s design is based on the RingMo. The architecture of the Swin Transformer consists of four stages, each comprising either a linear encoder or Patch Merging, along with an even number of Swin Transformer Blocks. Each block includes Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA). These attention mechanisms effectively compute attention maps within the windows, extracting spatial feature information between pixels while reducing the computational complexity of the model.
@@ 4.2 Model Architecture
To address the challenges above, we propose the overall architecture of RingMo-Aerial. As shown in Fig.7 Pre-training Stage 1, considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
@@
Figure 9: Positive samples for contrastive learning are obtained after the image’s affine transformation, cropping, and resizing.
Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module. To better understand the FE-MSA mechanism, it is essential first to recognize the limitations of the existing window attention mechanisms. These mechanisms often struggle to balance the semantic information in feature vectors of objects of different sizes. For instance, as illustrated in Fig.8, larger objects may occupy $4\times 4$ tokens, while smaller distant objects may occupy only $3\times 3$ tokens. The FE-MSA mechanism introduces frequency domain processing to the existing window attention, indirectly expanding the receptive field for smaller objects.
Fig.7(b) illustrates the RingMo-Aerial Block structure, including two window attention mechanisms optimized by the FE-MSA mechanism. Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block. Taking the first stage as an example, the input feature vector $z^{l}$ adopts the same slicing parameters as the standard Swin-Transformer, resulting in four-channel slices $C=\{c_{1},c_{2},c_{3},c_{4}\}$. After linear normalization, it enters the FE-MSA module. For each patch with a dimension of $C$, a $7\times 7$ convolution kernel is used for depthwise convolution to obtain the feature map $Fz^{(l-1)}=\text{Concat}\{Fc_{1},Fc_{2},Fc_{3},Fc_{4}\}$. Then, W-MSA or SW-MSA is used to compute the attention map. After the FE-MSA, small objects’ edge and texture features are further enhanced, improving the model’s ability to perceive local and global features. The computation process for each token patch through the RingMo-Aerial block is as follows:
@@
The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes. Fig.10 presents a comparative analysis of the feature extraction capabilities of the RingMo-Aerial and RingMo models in processing multiview ARS images. The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions. In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features. This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images. These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
4.3 Pre-training
Figure 10: Comparative analysis of feature heatmaps extracted from multi-view aerial imagery using RingMo-Aerial and RingMo. The feature heatmaps extracted by the RingMo-Aerial model are located in the upper half, while those from the RingMo model are in the lower half of the figure. In these images, the brighter the region, the more the model is focusing on that particular feature, indicating that these areas have higher feature importance for the model.
For pre-training, we construct a large-scale pre-training dataset. Our comprehensive pre-training dataset covers a wide range of aerial scenes, including high-altitude, low-altitude, and near-ground perspectives. It spans diverse geographical regions such as urban, rural, and wilderness areas. The dataset includes various targets, such as roads, vehicles, and pedestrians, and captures different seasons and periods. Notably, it exhibits multi-source, multi-temporal, and multi-instance characteristics. Before training, we crop the entire dataset into $640\times 640$ or $448\times 448$ resolution slices, all captured using optical three-channel imaging. The sliced dataset comprises approximately 5.4 million images (Fig.11).
Figure 11: The figure shows the number of slices corresponding to the slicing process performed on multiple collected open source airborne remote sensing datasets, which were used as the dataset for pre-training of the base model.
During the pre-training phase, we combine Masked Image Modeling (MIM) and Contrastive Learning (CL) techniques. For MIM training, we use the standard SimMIM pre-training framework. The primary goal of this approach is to enhance the model’s ability to capture pixel-level semantic associations, thus initially equipping the model with feature extraction capabilities. However, as analyzed previously, solely using the MIM pre-training method for ARS foundational model training is insufficient for the model to learn the intrinsic object characteristics specific to ARS vision. Therefore, building upon the MIM pre-training, we propose a contrastive learning framework based on affine transformations. This involves generating positive and negative image pairs through affine transformations, which serve as the basis for the contrastive learning pre-training of the model.
@@ 4.3 Pre-training
Among them, $\alpha$ is the transformation coefficient, and the value range is randomly between 0.05 and 0.35. On this basis, the perspective transformation matrix can be calculated. The perspective transformation matrix is defined as a $3\times 3$ matrix in the form of:
@@
The solution process of this transformation matrix is as follows:
@@
We obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$ through the aforementioned affine transformation process. The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs. By calculating the contrastive loss, we guide the model in learning the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
4.4 Downstream Fine-tuning
Figure 12: ARS-Adapter is an enhancement module designed to improve transformer models by adding efficient, low-cost adaptation layers. It integrates down-sampling, activation (e.g., GELU), and up-sampling, optionally including convolution, to enrich feature representations while reducing computational overhead.
TABLE I: Semantic Segmentation: Comp. on UAVID